OpenStudy (anonymous):

Aarush sat down to study at 8am the angle of elevation of the sun was 15 when he had finished studyng the angle of elevation of the sun was 55.For How long Did he Studied?

OpenStudy (anonymous):

he studied for (55-15)/360 * 24 hr=8/3 hr=2hr 40 min